OpenStudy (anonymous):

Graph y = -2x +3 and its inverse.

OpenStudy (anonymous):

answer choices

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@jigglypuff314 @jdoe0001

OpenStudy (jdoe0001):

if you recall, to find the "inverse relation" all we do is swap about the variables and solve for "y" so \(\bf {\color{brown}{ y}} = -2{\color{blue}{ x}} +3\qquad inverse\implies {\color{blue}{ x}}= -2{\color{brown}{ y}} +3\)

OpenStudy (jdoe0001):

then just graph each :)

OpenStudy (anonymous):

idk how to graph them

OpenStudy (anonymous):

is it C??

OpenStudy (jdoe0001):

ohh to graph them... just pick 2 random "x" and get a 'y" for each each one is just a line thus 2 points will make a line get the x,y for each random you picked and draw a line through them

OpenStudy (anonymous):

ok I tried but I got C am I correct

OpenStudy (jdoe0001):

hmmm what inverse function did you get?

OpenStudy (anonymous):

i just tried by using slope -2 and i thought C or D but how do i use your method

OpenStudy (jdoe0001):

well... parallel lines have the same slope, but that doesn't mean they have the same coordinates so... .to graph them... the slope alone won't do it

OpenStudy (anonymous):

oh ok but how do i do your method??

OpenStudy (jdoe0001):

the inverse? notice, all I did was swap the variables \(\bf {\color{brown}{ y}} = -2{\color{blue}{ x}} +3\qquad inverse\implies {\color{blue}{ x}}= -2{\color{brown}{ y}} +3\) then you need to solve for "y"

OpenStudy (anonymous):

ok how do i solve for y

OpenStudy (jdoe0001):

well... is just a linear equation.. I assume you've covered simplifying linears equations?

OpenStudy (anonymous):

yeah but i had a lot of trouble with that

OpenStudy (jdoe0001):

well for x=-2y+3 subtract 3 from both sides then divide both sides by -2 see what you get

OpenStudy (anonymous):

x-3/-2 = y

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

ok now what??

OpenStudy (jdoe0001):

then just pick a couple of random "x" let's try for say the function itself y = -2x+3 x=0 y = -2(0)+3 y=3 so the point will be (0,3) ----------------------------------------- let's pick another random "x" y= -2x+3 x=2 y=-2(2)+3 y=-4+3 y=-1 so the 2nd point for that will be (2, -1)

OpenStudy (jdoe0001):

so you have two points for the first one, and since the equation is just a line, you can just draw a line through them

OpenStudy (anonymous):

ok and then what

OpenStudy (anonymous):

I got A

OpenStudy (jdoe0001):

\(\bf x-3/-2 = y\\ \cfrac{x-3}{-2}=y\qquad x=0 \\ \quad \\ \cfrac{0-3}{-2}=y\implies \cfrac{3}{2}=y\\ \qquad \left(0,\cfrac{3}{2}\right)\\ ----------------------\\ \cfrac{x-3}{-2}=y\qquad x=2 \\ \quad \\ \cfrac{2-3}{-2}=y\implies \cfrac{1}{2}=y\\ \qquad \left(2,\cfrac{1}{2}\right)\)

OpenStudy (anonymous):

so was i correct A??

OpenStudy (jdoe0001):

ahemm recheck your graph

OpenStudy (anonymous):

D??

OpenStudy (jdoe0001):

check your graph

OpenStudy (anonymous):

im so confused is it the second graph??

OpenStudy (jdoe0001):

well... did you graph them? do they look like the 2nd graph?

OpenStudy (anonymous):

to me it looked like the last, what does yours look like

OpenStudy (anonymous):

thank you so much you helped me a lot :)

OpenStudy (jdoe0001):

yw